# Exploring the Potential of Encoder-free Architectures in 3D LMMs

### v2

@@ Abstract
Encoder-free architectures have been preliminarily explored in the 2D visual domain, yet it remains an open question whether they can be effectively applied to 3D understanding scenarios. In this paper, we present the first comprehensive investigation into the potential of encoder-free architectures to alleviate the challenges of encoder-based 3D Large Multimodal Models (LMMs). These challenges include the failure to adapt to varying point cloud resolutions and the point features from the encoder not meeting the semantic needs of Large Language Models (LLMs). We identify key aspects for 3D LMMs to remove the encoder and enable the LLM to assume the role of the 3D encoder: 1) We propose the LLM-embedded Semantic Encoding strategy in the pre-training stage, exploring the effects of various point cloud self-supervised losses. And we present the Hybrid Semantic Loss to extract high-level semantics. 2) We introduce the Hierarchical Geometry Aggregation strategy in the instruction tuning stage. This incorporates inductive bias into the LLM layers to focus on the local details of the point clouds. To the end, we present the first Encoder-free 3D LMM, Enel. Our 7B model rivals the current state-of-the-art model, ShapeLLM-13B, achieving 55.10%, 50.98%, and 43.10% on the classification, captioning, and VQA tasks, respectively. Our results demonstrate that the encoder-free architecture is highly promising for replacing encoder-based architectures in the field of 3D understanding. Code is released at https://github.com/Ivan-Tang-3D/ENEL.
Machine Learning, ICML
Figure 1: Issues of encoder-based 3D LMMs. (a) Point Cloud Resolution Limitation. During training, the point cloud size (P.T. size) and point token size (P.T. size) are fixed at 8192 and 512, respectively. And we adjust these two sizes during inference, point cloud size from 2K to 16K and the corresponding point token size from 128 to 2048. We evaluate them on the captioning task of the Objaverse benchmark using GPT-4 scores as the evaluation metric. (b) Embedding Semantic Discrepancy. We visualize the attention scores of the average text token to the point tokens, where red indicates higher values. The point tokens in the encoder-free architecture exhibit stronger textual semantic relevance needed for the LLM.
@@ 1 Introduction
Large Language Models (LLMs) (Touvron et al., 2023; Bai et al., 2023; Jiang et al., 2023; Cai et al., 2024) have gained unprecedented attention for their proficiency in understanding and generating complex language scenarios. Building upon these advances, many recent efforts have been made to develop Large Multimodal Models (LMMs), empowering LLMs with the capability to interpret multimodal information 2D images (Liu et al., 2024; Li et al., 2024) and 3D point clouds (Guo et al., 2023b; Xu et al., 2025), as well as generate multimodal asserts (Xie et al., 2024; Chen et al., 2025; Guo et al., 2025; Jiang et al., 2025).
@@
Mainstream LMMs are typically encoder-based, relying on heavyweight yet powerful multimodal encoders (e.g., CLIP (Radford et al., 2021) for 2D (Liu et al., 2021; Oquab et al., 2023) and I2P-MAE (Zhang et al., 2023a) for 3D). While these pre-trained encoders offer robust multimodal embeddings enriched with pre-existing knowledge, they also introduce challenges that could limit the future advancement of multimodal understanding.
@@
Specifically for 3D LMMs, the encoder-based architecture has the following potential drawbacks: (1) Point Cloud Resolution Limitation. 3D encoders are often pre-trained on point cloud data at a fixed resolution, such as 8,192 points for Point-BERT (Yu et al., 2022) in PointLLM (Xu et al., 2025). However, during inference, the resolution of point clouds may vary (e.g., 12,000 or 4,000 points). This difference between training and inference resolutions can result in the loss of spatial information when extracting 3D embeddings, leading to difficulties for LLMs to comprehend, as showcased in Figure 1 (a). (2) Embedding Semantic Discrepancy. 3D encoders are typically pre-trained using self-supervised methods like MAE (Pang et al., 2022; Tang et al., 2024a, b) and contrastive learning (Xie et al., 2020; Qi et al., 2023), but these training objectives may not align with the specific semantic needs of LLMs. In other words, they may not capture the most relevant semantics for LLMs to understand 3D objects, as visualized in Figure 1 (b). Even when a projection layer is used to connect 3D encoders with LLMs, simple MLPs are often insufficient for a complete semantic transformation. Given these issues, we ask: Is it possible to explore an encoder-free architecture for 3D LMMs, eliminating the 3D encoder and instead integrating its functionality directly within the LLM itself?
@@
In this paper, we present the first systematic investigation into the potential of an encoder-free architecture for 3D LMMs. To minimize external influences and ensure clarity, we use the pioneering and sufficiently concise PointLLM (Xu et al., 2025) as our encoder-based baseline, which consists of two progressive training stages: pre-training and instruction tuning. We evaluate the performance on 3D classification (Deitke et al., 2023) and 3D captioning (Deitke et al., 2023) tasks. Specifically, to remove the encoder while mitigating any performance degradation, we explore solutions to the following two key questions:
@@
(1) How can we compensate for the high-level 3D semantics originally extracted by the 3D encoder? In 3D LMMs, the raw point cloud input is first passed through a token embedding module for low-level tokenization, before being processed by the main 3D encoder, usually a Transformer (Vaswani, 2017), to generate high-level embeddings. Skipping the encoder entirely poses a challenge in capturing the complex spatial structures of 3D point clouds. To address this, we propose a strategy called LLM-embedded Semantic Encoding in the pre-training stage. First, we adopt a simple yet effective token embedding module that captures as much informative semantic content as possible. These 3D tokens are then directly fed into the LLM. Next, we aim to shift the responsibility of capturing high-level 3D semantics to the LLM itself. To facilitate this, we make the early layers of the LLM learnable, allowing them to specialize in multimodal alignment and further 3D encoding. To guide this process, we explore various 3D self-supervised loss functions, such as masked modeling loss and distillation loss, and ultimately propose the Hybrid Semantic Loss as the most effective choice.
Observation: Our adopted token embedding module, learnable LLM layers, and Hybrid Semantic Loss achieve comparable effectiveness to that of a pre-trained 3D encoder, effectively substituting it for high-level 3D semantics.
@@
(2) How can we integrate inductive bias into LLMs for better perception of 3D geometric structures? Traditional 3D encoders typically embed explicit inductive bias into their architectures to progressively capture multi-level 3D geometries. For instance, models like Point-M2AE (Zhang et al., 2022) use a local-to-global hierarchy, which is a concept also common in convolutional layers for 2D image processing (He et al., 2016). In contrast, LLMs employ standard Transformer architectures, where each layer processes the same number of tokens, representing the same semantic level across the network. In the absence of the encoder, we introduce the approach of Hierarchical Geometry Aggregation during the second fine-tuning stage. In the early layers of the LLM, we aggregate 3D tokens based on their geometric distribution using Farthest Point Sampling (FPS) and $k$-Nearest Neighbor ($k$-NN) sampling. This approach enables the LLM to gradually integrate detailed 3D semantics and develop a more holistic understanding of the 3D object. In the later layers, we reverse this aggregation, propagating the tokens back to their original distribution to maintain the fine-grained representation necessary for effective semantic communication.
Observation: We find that this hierarchical design can facilitate the acquisition of multi-level knowledge and better comprehend the 3D geometries of complex point clouds.
@@
Through a series of experimental investigations, we have uncovered the strong potential of applying encoder-free architecture to the 3D LMM domain. Building on our insights, we introduce Enel, an ENcoder-freE 3D LMM evolved from Vicuna-7B (Chiang et al., 2023) using the same training dataset from PointLLM. Notably, without any 3D encoders, Enel-7B achieves comparable performance to the current state-of-the-art ShapeLLM-13B (Qi et al., 2024), attaining scores of 55.10% and 50.98% on the classification and captioning tasks, respectively. We hope Enel may provide the community with a scalable and effective path for adapting the encoder-free architecture to 3D scenarios.
@@
• We further introduce Enel, a concise and well-performed encoder-free 3D LMM, which, at the 7B parameter scale, achieves 50.98%, 55.10% and 43.10% on 3D captioning, classification, and 3D VQA tasks, respectively, on par with existing encoder-based models.
Table 1: Token Embedding. Performance on Objaverse with PointLLM-7B as the baseline. ‘Cls’/‘Cap’: classification/captioning tasks. ‘Avg’: accuracy under prompts “What is this?” and “This is an object of.” ‘S-BERT’: Sentence-BERT. ‘T.E.’: our designed token embedding module.
Method Cls (Avg) Cap GPT-4 GPT-4 S-BERT PointLLM-7B 53.00 44.85 47.47 - Encoder 35.50 33.37 41.19 + 2-layer T.E. 42.50 41.35 44.25 + 3-layer T.E. 47.31 43.86 45.89 + 4-layer T.E. 45.00 42.99 44.51
@@
Method LR Cls (Avg) Cap GPT-4 GPT-4 S-BERT PointLLM-7B 2e-3 53.00 44.85 47.47 + 2 learnable layers 2e-3 41.06 42.23 45.92 4e-4 45.50 44.72 47.35 + 4 learnable layers 2e-3 44.85 41.53 46.77 4e-4 49.11 45.39 47.71 + 8 learnable layers 2e-3 43.76 39.71 42.38 4e-4 48.00 44.49 47.21 + 12 learnable layers 2e-3 42.91 40.05 42.40 4e-4 47.63 44.20 47.09
2 Related Work
3D LMM. Recent advancements in integrating large language models (LLMs) with 3D data have led to significant progress in both object-level and scene-level understanding. At the object level, early approaches like (Hong et al., 2024) utilize 2D rendering to leverage 2D LLMs, but this sacrifices geometric details. More recent models, including Point-Bind LLM (Guo et al., 2023b), PointLLM (Xu et al., 2023) and ShapeLLM (Qi et al., 2024), directly encode point clouds and align them with LLMs, by combining the 3D encoder with a powerful language model, effectively fusing geometric, appearance, and linguistic information. At the scene level, models like Chat-3D (Wang et al., 2023) and Scene-LLM (Fu et al., 2024) focus on understanding complex spatial relationships through dialogue and tasks like captioning. Scene-LLM (Fu et al., 2024) enhances embodied agents’ abilities in interactive 3D indoor environments by integrating both scene-level and egocentric 3D information. Grounded 3D-LLM (Chen et al., 2024b) utilizes referent tokens to reference specific objects within 3D scenes, enabling tasks such as object detection and language grounding.
Encoder-free Vision-Language Models. Traditional vision-language models (VLMs) often rely on vision encoders to extract visual features before processing them with language models, integrating image encoders like CLIP (Radford et al., 2021) and DINO V2 (Oquab et al., 2023). However, recent efforts have explored encoder-free VLMs for their simplicity. Approaches like (ChameleonTeam, 2024; Xie et al., 2024) use VQ tokenizers (Esser et al., 2021) or linear projection layers (Diao et al., 2024a; Chen et al., 2024a) to represent images. Fuyu-8B (Bavishi et al., 2023), a pure decoder-only model, directly processes image patches through linear projections, handling high-resolution images but showing only average performance. The EVE series (Diao et al., 2024b) eliminates the need for a separate vision encoder by bridging vision-language representation within a unified decoder and enhancing visual recognition capabilities through additional supervision.
@@ 3 Investigation of Encoder-free 3D LMM
Encoder-free modeling has been explored in the 2D vision domain to address issues related to image resolution and deployment overload. In this study, we conduct a comprehensive investigation to analyze the feasibility of adopting encoder-free architectures for 3D understanding tasks.
3.1 Overall Architecture
Task Formulation.
We present the first attempt to extend the encoder-free architecture to 3D LMMs, such as PointLLM (Xu et al., 2023) and ShapeLLM (Qi et al., 2024), in order to efficiently handle the complex tasks like embodied agent (Guo et al., 2023a). We select PointLLM as the baseline model for the exploration and evaluate the performance of different strategies on the Objaverse dataset (Deitke et al., 2023), using GPT-4 scores combined with traditional metrics as our evaluation metrics. The benchmark poses a significant challenge, requiring the model to align 3D semantics with text while capturing complex 3D geometries.
In the overall architecture, the encoder-free 3D LMM directly utilizes the token embedding module to convert point clouds into discrete point tokens, which are then concatenated with text tokens to serve as input to the LLM. As shown in Figure 2, to assume the role of the encoder, the LLM is guided to extract high-level semantic features of the point clouds and acquire multi-level knowledge from both global and local perspectives. In the subsequent sections, we primarily explore two strategies within the encoder-free architecture: LLM-embedded Semantic Encoding (Section 3.2) and Hierarchical Geometry Aggregation (Section 3.3).

### v4

@@ Abstract
Encoder-free architectures have been preliminarily explored in the 2D Large Multimodal Models (LMMs), yet it remains an open question whether they can be effectively applied to 3D understanding scenarios. In this paper, we present the first comprehensive investigation into the potential of encoder-free architectures to alleviate the challenges of encoder-based 3D LMMs. These long-standing challenges include the failure to adapt to varying point cloud resolutions during inference and the point features from the encoder not meeting the semantic needs of Large Language Models (LLMs). We identify key aspects for 3D LMMs to remove the pre-trained encoder and enable the LLM to assume the role of the 3D encoder: 1) We propose the LLM-embedded Semantic Encoding strategy in the pre-training stage, exploring the effects of various point cloud self-supervised losses. And we present the Hybrid Semantic Loss to extract high-level semantics. 2) We introduce the Hierarchical Geometry Aggregation strategy in the instruction tuning stage. This incorporates inductive bias into the LLM layers to focus on the local details of the point clouds. To the end, we present the first Encoder-free 3D LMM, Enel. Our 7B model rivals the state-of-the-art model, PointLLM-PiSA-13B, achieving 57.91%, 61.0%, and 55.20% on the classification, captioning, and VQA tasks, respectively. Our results show that the encoder-free architecture is highly promising for replacing encoder-based architectures in the field of 3D understanding. The code is released at https://github.com/Ivan-Tang-3D/ENEL.
@@ 1 Introduction
Large Language Models (LLMs) Touvron et al. (2023); Bai et al. (2023) have gained unprecedented attention for their proficiency in understanding and generating complex language scenarios. Building upon these advances, many recent efforts have been made to develop Large Multimodal Models (LMMs), empowering LLMs with the capability to interpret multimodal information, such as 2D images Li et al. (2024), 3D point clouds Chen et al. (2025); Guo et al. (2023); Xu et al. (2025); Wang et al. (2025) and visual generation Tong et al. (2025); Jiang et al. (2025); Guo et al. (2025b).
@@
Mainstream LMMs are typically encoder-based, relying on heavyweight yet powerful pre-trained encoders (e.g., CLIP Radford et al. (2021) for 2D and I2P-MAE Zhang et al. (2023a) for 3D). While these pre-trained encoders offer robust multimodal embeddings enriched with pre-existing knowledge, they also introduce challenges that could limit the future advancement of multimodal understanding. To mitigate the limitations introduced by visual encoders in VLMs—such as resolution, aspect ratio, and semantic priors—many encoder-free LMM studies Li et al. (2025); Diao et al. (2024a; 2025); Lei et al. (2025); Luo et al. (2025) have explored the possibility of training without pre-trained encoders.
Figure 1: Issues of encoder-based 3D LMMs. (a) Point Cloud Resolution Limitation. During training, the point cloud size (P.C. Size) and point token size (P.T. Size) are fixed at 8192 and 512, respectively. And we adjust these two sizes during inference, point cloud size from 2K to 16K and the corresponding point token size from 128 to 2048. We evaluate them on the captioning task of the Objaverse benchmark using GPT-4 score as the evaluation metric. (b) Embedding Semantic Discrepancy. We visualize the attention scores of the average text token to the point tokens, where red indicates higher values. The point tokens in the encoder-free architecture exhibit stronger textual semantic relevance needed for the LLM.
@@
Specifically for 3D LMMs, the encoder-based architecture has the following potential drawbacks: (1) Point Cloud Resolution Limitation. 3D encoders are often pre-trained on point cloud data at a fixed resolution, such as 8,192 points for Point-BERT Yu et al. (2022) in PointLLM Xu et al. (2025). However, during inference, the resolution of point clouds may vary (e.g., 12,000 or 4,000 points). This difference between training and inference resolutions can result in the loss of spatial information when extracting 3D embeddings, leading to difficulties for LLMs to comprehend, as showcased in Figure 1 (a). (2) Embedding Semantic Discrepancy. 3D encoders are typically pre-trained using self-supervised methods like MAE Pang et al. (2022); Tang et al. (2024a; b) and contrastive learning Xie et al. (2020); Qi et al. (2023), but these training objectives may not align with the specific semantic needs of LLMs. In other words, they may not capture the most relevant semantics for LLMs to understand 3D objects, as visualized in Figure 1 (b). Even when a projection layer is used to connect 3D encoders with LLMs, simple MLPs are often insufficient for a complete semantic transformation. Given these issues, we ask: Is it possible to explore an encoder-free architecture for 3D LMMs, eliminating the 3D encoder and instead integrating its functionality directly within the LLM itself?
@@
In this paper, we present the first systematic investigation into the potential of an encoder-free architecture for 3D LMMs. To minimize external influences and ensure clarity, we use the pioneering and sufficiently concise PointLLM Xu et al. (2025) as our encoder-based baseline, which consists of two progressive training stages: pre-training and instruction tuning. We evaluate the performance on 3D classification Deitke et al. (2023), 3D captioning Deitke et al. (2023) and 3D VQA Deitke et al. (2023) tasks. Specifically, to remove the encoder while mitigating any performance degradation, we explore solutions to the following two key questions:
@@
(1) How can we compensate for the high-level 3D semantics originally extracted by the 3D encoder? In 3D LMMs, the raw point cloud input is first passed through a token embedding module for low-level tokenization, before being processed by the main 3D encoder, usually a Transformer Vaswani (2017), to generate high-level embeddings. Skipping the encoder entirely poses a challenge in capturing the complex spatial structures of 3D point clouds. To address this, we propose a strategy called LLM-embedded Semantic Encoding in the pre-training stage. First, we adopt a simple yet effective token embedding module that captures as much informative semantic content as possible. These 3D tokens are then directly fed into the LLM. Next, we aim to shift the responsibility of capturing high-level 3D semantics to the LLM itself. To guide this process, we explore various 3D self-supervised loss functions, such as masked modeling loss and distillation loss, and ultimately propose the Hybrid Semantic Loss as the most effective choice. Further, we make the early layers of the LLM to be learnable, allowing them to specialize in multimodal alignment.
@@
(2) How can we integrate inductive bias into LLMs for better perception of 3D geometric structures? Pre-trained 3D encoders typically embed explicit inductive bias into their architectures to progressively capture multi-level 3D geometries. For instance, models like Point-M2AE Zhang et al. (2022) use a local-to-global hierarchy, which is a concept also common in convolutional layers for 2D image processing He et al. (2016). In contrast, LLMs employ standard Transformer architectures, where each layer processes the same number of tokens, representing the same semantic level across the network. In the absence of the encoder, we introduce the approach of Hierarchical Geometry Aggregation during the fine-tuning stage. In the early layers of the LLM, we aggregate 3D tokens based on their geometric distribution using Dynamic Grid Sampling. This approach enables the LLM to gradually integrate detailed 3D semantics and develop a more holistic understanding of the 3D object. In the later layers, we reverse this aggregation, propagating the tokens back to their original distribution to maintain the fine-grained representation necessary for complex tasks.
@@
Through a series of experimental investigations, we have uncovered the strong potential of applying encoder-free architecture to the 3D LMM domain. Building on our insights, we introduce Enel, an ENcoder-freE 3D LMM evolved from Vicuna-7B Chiang et al. (2023) using the same training dataset from PointLLM. Notably, without any 3D encoders, Enel-7B achieves comparable performance to the state-of-the-art PointLLM-PiSA-13B Guo et al. (2025a). We hope Enel may provide the community with an effective path for adapting the encoder-free architecture to 3D scenarios.
@@
• We further introduce Enel, a concise and well-performed encoder-free 3D LMM, which, at the 7B parameter scale, achieves 57.91%, 61.0%, and 55.20% on 3D captioning, classification, and 3D VQA tasks, respectively, on par with existing encoder-based models.
2 Related Work
3D LMM. Recent advancements in integrating large language models (LLMs) with 3D data have led to significant progress in both object-level and scene-level understanding. At the object level, early approaches like Hong et al. (2024) utilize 2D rendering to leverage 2D LLMs, but this sacrifices geometric details. More recent models, including Point-Bind LLM Guo et al. (2023), PointLLM Xu et al. (2023b) and ShapeLLM Qi et al. (2024), directly encode point clouds and align them with LLMs, by combining the 3D encoder with a powerful language model, effectively fusing geometric, appearance, and linguistic information. MiniGPT-3D Tang et al. (2024c) is introduced, which efficiently aligns 3D point clouds with LLMs by leveraging 2D priors from 2D-LLMs. It employes a four-stage cascaded training strategy along with a Mixture of Query Experts (MoQE) module. Zeng et al. propose GreenPLM Tang et al. (2025), an energy-efficient framework that directly translates monolingual pre-trained language models into other languages using bilingual lexicons. At the scene level, models like Chat-3D Wang et al. (2023) and Scene-LLM Fu et al. (2024) focus on understanding complex spatial relationships through dialogue and tasks like captioning. Scene-LLM Fu et al. (2024) enhances embodied agents’ abilities in interactive 3D indoor environments by integrating both scene-level and egocentric 3D information. Grounded 3D-LLM Chen et al. (2024b) utilizes referent tokens to reference specific objects within 3D scenes, enabling tasks such as object detection and language grounding. However, conventional encoder-based 3D LMMs commonly suffer from limitations, specifically Point Cloud Resolution Limitation and Embedding Semantic Discrepancy, which stem from the inductive bias inherent in the 3D pre-trained encoder. Our Enel alleviates these restrictions by removing the encoder and utilizes a lightweight architecture to significantly boost performance.
Encoder-free Vision-Language Models. Traditional vision-language models (VLMs) often rely on vision encoders to extract visual features before processing them with language models, integrating image encoders like CLIP Radford et al. (2021) and DINO V2 Oquab et al. (2023). However, recent efforts have explored encoder-free VLMs for their simplicity. Approaches like ChameleonTeam (2024); Xie et al. (2024) use VQ tokenizers Esser et al. (2021) or linear projection layers Diao et al. (2024a); Chen et al. (2024a) to represent images. Fuyu-8B Bavishi et al. (2023), a pure decoder-only model, directly processes image patches through linear projections, handling high-resolution images but showing only average performance. The EVE series Diao et al. (2024b; 2025) eliminates the need for a separate vision encoder by bridging vision-language representation within a unified decoder and enhancing visual recognition capabilities through additional supervision. Mono-InternVL series Luo et al. (2024; 2025) leverage visual experts and progressive visual pre-training (EViP/EViP++) to achieve stable optimization and competitive performance. SAIL series Lei et al. (2025) directly encode raw pixels and decodes language within a single architecture, achieving competitive vision-language performance without pre-trained vision encoders. The key idea behind Enel is enabling the LLM to assume the functionality of the encoder by effective and efficient methods. This approach diverges from 2D encoder-free LMMs, which tend to focus on larger datasets and more complex structures for better results.
Table 1: Token Embedding. Performance on Objaverse with PointLLM-7B as the baseline. ‘Cls’/‘Cap’: classification/captioning tasks. ‘Avg’: accuracy under prompts “What is this?" and “This is an object of." ‘S-BERT’: Sentence-BERT. ‘T.E.’: our designed token embedding module.
Method Cls (Avg) Cap GPT-4 GPT-4 S-BERT PointLLM-7B 53.00 44.85 47.47 - Encoder 35.50 33.37 41.19 + 2-layer T.E. 40.60 38.85 43.25 + 3-layer T.E. 45.55 41.36 44.82 + 4-layer T.E. 43.00 40.47 43.50
@@ 2 Related Work
Method LR Cls (Avg) Cap GPT-4 GPT-4 S-BERT PointLLM-7B 2e-3 53.00 44.85 47.47 + 2 learnable layers 2e-3 40.00 40.20 44.82 4e-4 44.00 42.62 46.30 + 4 learnable layers 2e-3 43.75 40.13 45.76 4e-4 47.90 43.50 46.70 + 8 learnable layers 2e-3 42.35 37.91 41.28 4e-4 46.70 42.80 46.14 + 12 learnable layers 2e-3 41.55 40.05 41.40 4e-4 46.15 42.39 46.00
@@ 3 Investigation of Encoder-free 3D LMM
3.1 Preliminary
Encoder-free in 2D LMMs. ELVA Li et al. (2025) is an encoder-free Video-LLM that directly models nuanced video-language interactions without relying on a vision encoder. EVE Diao et al. (2024a) and its successor EVEv2 Diao et al. (2025) are designed as efficient encoder-free vision-language models. SAIL Lei et al. (2025) serves as a unified transformer for vision and language, while Mono-InternVL Luo et al. (2025) represents a monolithic multimodal LLM. In parallel, Fuyu-8B Bavishi et al. (2023), a decoder-only transformer developed by Adept AI, has gained substantial community adoption. A common characteristic across these works is the adoption of a lightweight, randomly initialized token embedding layer to convert inputs into tokens for the LLM. This design eliminates the need for a dedicated vision encoder and enables end-to-end training and inference.
Pre-trained Encoders in 3D LMMs. Traditionally, 3D pre-trained encoders are characterized by two properties: (1) independent pretraining on point cloud tasks (e.g., reconstruction), and (2) structural decoupling, where they are connected to the language model through projection layers. In 3D LMMs, commonly adopted encoders refer to pre-trained models such as PointMAE Pang et al. (2022), PointBERT Yu et al. (2022), and Uni3D Zhou et al. (2023). Related work is in Appendix A.1.
Overall Architecture. We select PointLLM as the baseline model for the exploration and evaluate the performance of different strategies on the Objaverse dataset Deitke et al. (2023), using GPT-4 scores combined with traditional metrics as our evaluation metrics. Point Embedding Layer. As shown in Figure 2, we first remove the encoder of PointLLM and adopt the original token embedding Yu et al. (2022). However, the coarse structural design results in a significant performance degradation, as observed in Table 2, where the GPT-4 scores for the classification and captioning tasks decrease by 17.5% and 10.48%, respectively. To mitigate excessive information loss and provide refined local features to the LLM, we adopt a small network with a limited number of parameters, which is a lightweight variant of Point-PN Zhang et al. (2023b). Specifically, for the input $\{P_{i}\}_{i=1}^{N}$, we apply Farthest Point Sampling (FPS) for downsampling the number of points, k-Nearest Neighbors (k-NN) with group size k for local aggregation, and learnable linear layers for feature encoding. After a series of repetitive operations and the projection layer, we transform the point clouds into high-dimensional vectors $\{F_{i}\}_{i=1}^{M}\in\mathbb{R}^{M\times D_{1}}$. In Table 2, we experiment with token embedding at different depths and find that three layers yield the best performance. 3D Encoding & Alignment. We discover that the absence of the encoder results in a lack of context modeling in point cloud feature processing. Therefore, we attempt to have the early layers of the LLM take on the encoder’s role in capturing global interactions of features, further encoding the point cloud features. In the pre-training stage, we set the first K layers of the frozen LLM to be learnable. Within the shared semantic space, 3D tokens and text tokens interact and align naturally. Early Fusion provides a more practical way to achieve modality alignment between 3D and textual semantic spaces. Meanwhile, we experiment with different learning rates. As shown in Table 2, a smaller learning rate yields better results by stabilizing early layer optimization. Based on the designed token embedding module, setting the first four layers to be learnable yields the best results.
3.2 LLM-embedded Semantic Encoding
The lack of the 3D encoder results in insufficient encoding of point cloud semantic information, which greatly hinders the LLM to understand the structural details of point clouds. Most existing 3D encoders use self-supervised losses to embed the high-level semantics of point clouds into the transformer, primarily categorized into four types: Masked Modeling Loss Pang et al. (2022), Reconstruction Loss Qi et al. (2023), Contrastive Loss Khosla et al. (2020), and Knowledge Distillation Loss Zhang et al. (2023a). Based on the proposed token embedding module and LLM learnable early layers, we implement and evaluate the effects of these losses on the encoder-free 3D LMM in the pre-training stage, as described in Figure 3. Finally, we propose the Hybrid Semantic Loss, which assists the LLM to learn the relationship between local spatial information in the point clouds and grasp the high-level 3D semantics.
Masked Modeling Loss. In the pre-training stage, we apply the Masked Modeling Loss to the point tokens processed by the LLM, as shown in Figure 3 (a). Through the token embedding module, the point clouds $\{P_{i}\}_{i=1}^{N}$ are divided into point patches $\{G_{i}\}_{i=1}^{M}\in\mathbb{R}^{M\times k\times 3}$ and the corresponding point tokens $\{F_{i}\}_{i=1}^{M}$. We randomly mask the point tokens with a masking ratio $r$, and replace them with learnable tokens. The masked feature tokens can be denoted as $\{F_{\text{gt}_{i}}\}_{i=1}^{M*r}$, which serve as the ground truth for the loss computation. After the masked tokens are replaced with learnable tokens and processed by the LLM, a linear layer is applied to predict the point tokens $\{F_{\text{pre}_{i}}\}_{i=1}^{M*r}\in\mathbb{R}^{M*r\times D_{1}}$, and the Mean Squared Error (MSE) is computed between $F_{\text{pre}}$ and $F_{\text{gt}}$. The optimization is:
The specific process of applying Masked Modeling to point patches $G$ is detailed in Appendix A.3.1.
where $a=G_{\text{pre}}$, $b=G$. The procedure for reconstructing feature F is detailed in Appendix A.3.1.